Abey Ltd entered into a contract with Data Pro Ltd for the lease of twenty-four laptop computers. The contract was entered into on 1 April 20x1 for a 2-year period. Each item is of low value and Abey Ltd applies the low value exemption of IFRS 16 Leases. The benefit derived for Abey Ltd from the lease agreement is constant over the lease period.

The following amounts are payable to Data Pro Ltd per the lease agreement:

• From 1/4/20x1-31/03/20x2  $2 000 per month

• From 1/4/20x2-31/03/20x3​$3 000 per month

Abbey Ltd has a 31 December year-end

Required:

Prepare the journal entries for the accounting records of Abey Ltd for the year ended 31 December 20x1, 20x2 and 20x3.



1
Expert's answer
2021-06-14T13:15:29-0400

Solution:

Lease payments for the year ended December 20X1 = 9 months "\times" 2,000 = 18,000

Lease payments for the year ended December 20X2 = (3 months "\times" 2,000) + (9 months "\times" 3,000) = 6,000 + 27,000 = 33,000

Lease payments for the year ended December 20X1 = 3 months "\times" 3,000 = 9,000

 

The journal entries will be entered as follows:

 

                                                                                                            Dr.                          Cr.

December 20X1                      Lease liability                                    18,000

                                                           Cash                                                                        18,000

(To record lease payments for the year ended December 20X1)

 

                                                                                                            Dr.                          Cr.

December 20X2                      Lease liability                                   33,000

                                                           Cash                                                                        33,000

(To record lease payments for the year ended December 20X2)

 

                                                                                                            Dr.                          Cr.

December 20X3                      Lease liability                                   9,000

                                                           Cash                                                                        9,000

(To record lease payments for the year ended December 20X3)